• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:02:03 on December 10, 2007

    Entry number 219702

    keyword=Shift Summary

    Day Shift 12/10/07
    John LeRose (SL)
    Konrad Aniol (TO)
    Jose Udias (3rd)

    Ebeam=739.4349 MeV (HALLA:p)
    Left: P0=0.1828 90°
    Right: P0=0.14964 (settling in) 90°
    start run# Target
    08:04 3656 LH2
    08:05 3657 15 cm dummy
    08:16 3658 Carbon
    08:33 3659 Fe

    08:45 call MCC for Controlled Access for left arm polarity change
    negative to positive. Techs notified
    09:20 out of access called for beam

    Left +: P0=0.14964 90°
    Right-: P0=0.14964 90°

    09:39 3660/21708 Fe (40 microamps)
    09:58 3661/21709 Carbon (40 microamps)
    10:15 21710 15 cm LH2 (40 microamps)
    10:16 3662/21711 15 cm LH2 (40 microamps)
    10:29 3663/21712 15 cm LH2 + Pb (10 microamps)

    Left +: P0=0.13539 90° (change)
    Right-: P0=0.14964 90° (no change)

    10:51 3664 15 cm LH2 + Pb (10 microamps)
    11:06 3665/21713 15 cm LH2 (40 microamps)
    11:32 3666/21714 Carbon (40 microamps)
    11:45 3667/21715 Fe (40 microamps)

    Left +: P0=0.12249 90° (change)
    Right-: P0=0.14964 90° (no change)

    11:57 21716 Fe (40 microamps)
    12:03 3668 Fe (40 microamps)
    12:05 3669 Fe (40 microamps)
    12:20 3670/21717 carbon (40 microamps)
    12:39 3671/21718 15 cm LH2 (40 microamps)
    12:49 3672/21719 15 cm LH2 + Pb (10 microamps)

    Left +: P0=0.11083 90° (change)
    Right-: P0=0.14964 90° (no change)

    13:09 3673/21720 15 cm LH2 + Pb (10 microamps)
    13:26 3674/21721 15 cm LH2 (40 microamps)
    13:39 21722 carbon (40 microamps)
    13:41 3675 carbon (40 microamps)
    13:42 3676 carbon (40 microamps)
    13:44 beam off for a "couple of minutes"
    13:49 3677/21723 carbon (40 microamps)
    14:02 3678/21724 Fe (40 microamps)

    14:12 Finished positron running called for controlled access
    for left arm polarity change and spectrometer move to 120° (both arms)

    Accelerator planning to do beam tuning for Hall C
    Set right arm to 0.71197 GeV/c (Q2/3 cycled)

    At shift end:
    Spectrometers in motion
    Right arm P0 is set (quads cycled)
    Left dipole ramping, quads not set (need to be cycled)